# HandProxy: Expanding the Affordances of Speech Interfaces in Immersive Environments with a Virtual Proxy Hand

### v1

@@ 5.3. Data Collection and Analysis
Figure 7. Command diversity in test sessions: Each task was repeated on three objects — one for practice and two for test sessions (top row: second object, bottom row: third object). We report the number of unique command phrasings (#), execution accuracy (Rate), normalized entropy ($H_{N}$), and a histogram showing the distribution of unique phrasings. Finally, we provide examples of successful and unsuccessful participant commands.
[…]
@@ 6.1. RQ1: To what extent can participants complete a wide variety of tasks using HandProxy?
Participants successfully completed a wide variety of tasks using HandProxy, with an 100% task completion rate on tasks conducted in the test sessions. Among the 781 commands participants issued across all test sessions, the system correctly executed 717 of them, resulting in an overall accuracy of 91.8%. The median command execution accuracy per participant is 92.5%, with an IQR of 89.3% to 95.5%. On average, the system took 1.66 seconds to interpret the recognized speech command, with an standard deviation of 0.94 seconds.
After errors happened, participants were able to recover from errors by using alternative commands to complete the task. Multiple techniques were reported for error recovery, including rephrasing (13), repeating (3), and splitting into smaller steps (3). Across all commands in the test sessions, an average of 1.09 attempts were needed for a command to succeed. Specifically, 85.5% of error commands were resolved with 1 more attempt, 12.7% took 2 more attempts, and 1.8% took 3 more attempts, showing that most errors were corrected by giving one alternative attempt.
@@ 6.2. RQ2: How effectively can the system interpret the diverse ways in which users issue commands?
HandProxy was able to handle users’ diverse ways of phrasing commands effectively, including those less frequently used. To identify unique commands per task, we applied lemmatization and removed stopwords from each command used by participants during test sessions. Figure 7 shows the histogram and normalized entropy of unique commands participants used to complete study tasks. In the histograms, failed attempts are highlighted in red, while successful attempts are shown in blue. The normalized entropy quantifies the variation in the distribution of commands, with values ranging from 0 to 1. A value of 0 indicates that all attempts were dominated by a single command, while a value of 1 shows that each attempt corresponds to a unique command. Both the histogram and normalized entropy shows the diversity of commands participants used. Specifically, tasks were not dominated by just one command but were distributed across various commands, and the system was able to interpret most commands, including those used less frequently.
@@
During the practice sessions, participants generated an average of 32 unique commands per task. The system successfully interpreted variations of commands, including different verbs (e.g., {grab, pick up, fetch, hold} the peach), object descriptions (e.g., peach, pink fruit), spatial references (e.g., first fruit, second watermelon from the left, the watermelon in the middle), and sentence structures (e.g., can you use the minimize button, peach inside the basket). Moreover, the system demonstrated the ability to interpret and decompose high-level commands and generalize beyond literal meanings. For example, it interpreted “increase the brightness” as (i) grab the brightness knob and (ii) twist right. In the example of “fold the window,” the system was able to connect fold to minimize, and press the minimize button.
@@
After practice sessions, as shown in Figure 7, participants continued to use diverse commands and phrasings rather than relying solely on the most straightforward command, and the system effectively handled this variability. Given that participants were free to use their own words without predefined commands or formats, these results highlight HandProxy’s capability to interpret flexible and varied input.
@@
To better understand HandProxy’s limitation, we analyzed the retried commands in the test sessions. In total, the system failed to execute or incorrectly executed 64 commands. Among them, 40 were categorized as invalid, and the system prompted users to try again and did not perform any action. Most were due to challenges in identifying target objects (n=26), including synonyms (9), visual descriptions (6), object functionality (6), positional constraints (3), or other system errors (2). For example, participants described the power button as “the white and round button on the right side of the screen” (P1), the basket as “the brown object” (P16), or directly described the visual content within a widget, such as describing the window as “the dog photo” (P2). In other cases, participants referred to objects with synonyms or their expected functionality, such as describing the power button as “start on off button” (P10), or the heart button as “favorite button” (P16). While in many cases the system could infer beyond the object metadata, it struggled to account for the possible descriptions described above. This highlights the need for a more comprehensive understanding of the objects and environment to support more diverse descriptive commands, which we will discuss in future work.
@@
Apart from invalid commands with no action, the system incorrectly executed 24 valid commands. Among these, 3 were due to speech recognition errors, while 18 resulted from misinterpreted hand interactions. These include ambiguous commands (6), incorrect command decomposition (5), incorrect gesture parsing (4), or performed on the wrong target objects (3). Some commands had multiple possible interpretations, where context could have clarified the intended meaning. For instance, when P6 said, “turn the volume slider to maximum,” the participant expected the slider to move to the top. However, the system interpreted the word “turn” literally and performed a “twist right” gesture instead. Errors also occurred when participants combined multiple commands into one sentence. For example, P20 said, “pinch the resize button and pull it to the right” to make a window wider, expecting a continuous pinch-and-move gesture. Instead, the system parsed “pull” as a separate command, initiated a “grab” gesture that canceled the pinch and caused an unexpected action. The system identified the wrong object from some other commands. For example, P5 said “pick up the first watermelon on the right” to refer to the rightmost watermelon, but the system interpreted it as the first watermelon overall (left to right).
@@
Nonetheless, participants were able to complete 100% of the tasks and recover from errors using alternative commands, with an average of 1.09 attempts per command.
@@ 6.3. RQ3: What strategies, preferences, and methods do participants use for specifying intent?
Participants demonstrated diverse strategies and preferences while prompting the system. During practice sessions, the average command length was 5.25 words (median: 5 words). In test sessions, this was slightly lower, with an average of 4.73 words and a median of 4. For free exploration sessions—where users were not restricted to specific tasks—the average command length increased to 5.95 words, with a median of 5. Additionally, the command length also exhibits a large variance, which indicates participants’ diverse approaches to construct commands. Some participants preferred longer, more detailed commands, as they found it easier to “specify what I want” (P2) or believed it would “help the system understand” their intent more accurately (P7). For example, to grab the peach, P15 used the command, “can you grab the peach and hold it in your hand?” On the other hand, many participants still preferred keeping the commands short. In the interviews, 14 out of 20 participants favored shorter commands for their simplicity, clarity, and lower risk of errors. As P16 mentioned, “The longer it is, the more words I could say wrong and it could misinterpret.”
However, having shorter commands does not necessarily mean commands are always simple and low-level. In fact, participants used a mix of high-level and direct commands, even when the command is short. For example, P9 used “maximize the volume” to let the system drag the volume to the top. At the same time, shorter commands could be a consequence of decomposing task into smaller steps. For the same tasks mentioned above, P12 said “pinch the volume slider” followed by “pull up”. This is also reflected in interviews, where 9 participants preferred high-level commands, 5 preferred detailed controls, and 6 indicated that it depends on the specific task. For example, P9 mentioned that “I just wanted to tell the system what I want to do, and let the system figure out what gestures to do,” while others mentioned that using detailed control would be “more accurate” (P8), especially in precise tasks (P11) or those that need fine grained controls (P6). This demonstrated the necessity of supporting commands of different levels in hand control, which highlights the benefits HandProxy could bring compared to limited command-based voice controls.
To get a deeper understanding of how different users would prompt the system, we asked participants to discuss their prompting strategies during interviews. We found that at first, some participants needed to switch to a new mental model from doing to describing the hand interactions. As P17 mentioned:
@@
This led some participants to use more descriptive ways of specifying their interaction intentions. For example, P13 used “touch your index finger and thumb” to describe the pinch gesture in the warm-up tasks, and P5 described the whole motion of pick up peach, move up, left, release hand to finish the put the peach into the basket task in the practice session. This descriptive behavior also applied to objects as well. As described in subsection 6.1, participants used different descriptive ways to identify the target object. Utilizing the common sense encapsulated in LLMs, HandProxy was able to correctly interpret these variations in cases where users were unsure about how a specific hand motion or object should be described.
@@ 6.4. RQ4: How do participants experience and perceive HandProxy?
We evaluated system usability using 7-point Likert scale questions, covering system effectiveness, interaction variety, ease of use, responsiveness, learnability, consistency, and user confidence. Overall, participants found HandProxy effective for hand interactions (avg. 5.5, std. 0.82), can support various interaction gestures (average 5.75, std. 0.91), easy to use (average 5.7, std. 1.12), and participants were generally confident in using it (average 5.3, std. 1.08). Regarding system responsiveness, participants gave an average of 5.15 (std. 1.18). While participants appreciated the fast timing controls (e.g., stop, continue, undo) (e.g., P6), overall responsiveness across different types of commands could be improved for smoother interaction (P2, P8). HandProxy was also seen as intuitive and easy to learn (avg. 5.7, std. 1.41), especially given that the study was intentionally designed to have only a very brief on-boarding session. As P3 mentioned, “a few minutes of play and you should be ready.” However, P15 highlighted challenges for users unfamiliar with AR/VR, often relying on long, descriptive commands and struggling with terminology. Both P15 and P11 suggested that a tutorial with example commands, words, and objects would make the system more intuitive for beginners.
During the interviews, participants highlighted the system’s effectiveness in performing various hand interactions and its intuitiveness to use. Specifically, participants valued its ability to handle high-level commands (P1, P2, P8, P11, P19); robustness to certain ambiguity in the command (P5, P10, P11); consistency on the same task (P7, P15); its memory capability (P19); and its ability to take continuous speech input and perform actions on the go (P7). For example, P1 mentioned that HandProxy is able to directly execute her high-level commands step by step, “even though I haven’t told the system how to perform the actual task.” P7 mentioned the efficiency the continuous speech input could bring, and said “Another thing it did pretty well is, if you gave it a complex command of doing multiple things, it would do those things without any hesitation between them. And once I figured out you could do that, it was a lot faster than doing this like one command and then waiting for it to happen before you give it the next one.”
Figure 8. Likert scale responses of system usability.
@@
A figure of Likert scale responses across seven evaluation criteria related to the system’s usability and performance. The categories on the vertical axis include: Effectiveness of Initiating Hand Interactions, Variety of Interactions User Can Perform, Easiness of Use, Responsiveness of the System, Easiness to Learn, Consistency of the System Performance, Confidence of Using the System for Interaction Tasks. The majority of responses lean towards agreement and strong agreement, particularly in categories of ”Easiness of Use,” ”Easiness to learn,” ”System consistency,” and ”System variety”
@@ 6.5. RQ5: What other expectations do participants have for the system?
To gather additional expectations, here we report patterns among commands participants tried in the free exploration sessions, and summarize suggestions participants gave during the interview. In the free exploration session, many participants attempted commands that required multimodal understanding of the environment, context, and how different objects are related to each other. For example, P18 wanted to say “like the dog” to press the heart button on top of the image viewer window. This requires the system to not only have the knowledge that the image viewer is currently showing the dog, but also to spatially and semantically connect the heart button to the dog. Many other examples also relied on environment understanding, such as visual features (“grab the 3rd green object” by P18) and spatial relationships (“grab the watermelon between the two fruits” by P16). This could not only provide more flexibility in speech commands, but also help identify the object users do not know how to describe. As P19 mentioned,
@@
The expectation of environment understanding also extends to how the physics work in the virtual environment, such as “throw the apple to the dog,” “catch the apple right before it hits the table,” and “move the apple from the table to the floor.” The expected actions and the movement path usually depend on how the virtual world works, such as the hitbox of the table so that the hand can go around, rather than go through, the table. Since the current system does not support complex environment understanding beyond text-based metadata, supporting real-time multimodal understanding could enhance its expressiveness and flexibility.
Participants also tried additional gestures that the system did not support, such as roll, throw, flip, wipe, and detailed finger controls. While there was no direct match to these gestures, HandProxy still tried to utilize existing gestures to reproduce the command to its best. For example, the system performed a grab and rotate to “flip a basket,” and swipe left and right multiple times when asked to “clean the table.” In addition, participants also suggested other hand controls that could be useful, including bimanual control (P1, P2, P13, such as pushing using two hands, grabbing different objects with each hand), supporting more precise object manipulation (P2, such as grabbing the handle of the basket versus the side of the basket), and detailed finger controls (P5, such as bending the first two fingers). HandProxy is designed to be extensible, and these hand interactions could be integrated in the future.
Additionally, participants would like the system to understand object states and conditions in the environment. For example, P20 tried “if the photo has been resized, press the confirm button,” and P6 tried “increase the volume to the same level as the brightness.” This also includes monitoring the state of the object while being manipulated by the virtual hand. For example, P13 mentioned that, “There was one time when it was trying to move the apple but didn’t realize it no longer had the apple in its hand,” if the system was aware of the hand state, it could automatically correct this error. These require understanding and monitoring the object state to relieve users from “monitoring conditions constantly.” (P19)
Lastly, although the current system displays recognized results and previews the hand movement path, participants emphasized the importance of additional feedback to help users identify and recover from errors. For example, P6 recommended underlining the parts in user’s command that are confusing or uninterpretable, enabling users to rephrase their input more effectively in subsequent attempts. P1 suggested having the system predict and suggest possible next moves after the command, such as showing all afforded hand interactions once the user grabbed an object. Additionally, participants noted that improving the speech recognition system and reducing overall latency could enhance the system’s usability. These suggestions highlight key areas for future development to further improve the system’s functionality and user experience.

### v2

@@ 6.1. RQ1: To what extent can participants complete a wide variety of tasks using HandProxy?
Participants successfully completed various tasks using HandProxy, with an 100% task completion rate on tasks conducted in the test sessions. Among the 781 commands participants issued across all test sessions, the system correctly executed 717 of them, with an overall accuracy of 91.8%. The median command execution accuracy per participant is 92.5%, with an IQR of 89.3% to 95.5%. On average, the system took 1.66 seconds to interpret the recognized speech command, with an standard deviation of 0.94 seconds.
After errors happened, participants were able to recover from errors by using alternative commands to complete the task. Multiple techniques were reported for error recovery, including rephrasing (13), repeating (3), and splitting into smaller steps (3). Across all commands in the test sessions, an average of 1.09 attempts (std 0.33) were needed for a command to succeed. Specifically, 85.5% of error commands were resolved with 1 more attempt, 12.7% took 2 more attempts, and 1.8% took 3 more attempts, showing that most errors were corrected with one alternative attempt.
@@ 6.2. RQ2: How effectively can the system interpret the diverse ways in which users issue commands?
HandProxy effectively handled users’ diverse ways of phrasing commands, including less common ones. To identify unique commands per task, we lemmatized and removed stopwords from commands used by participants during test sessions. Figure 7 shows the histogram and normalized entropy of unique commands participants used to complete study tasks. The normalized entropy quantifies the variation in the distribution of commands with a value from 0 (all attempts were the command) to 1 (each attempt corresponds to a unique command). Both the histogram and normalized entropy shows the diversity of commands participants used. Specifically, tasks were not dominated by just one command but were distributed across various commands, and the system was able to interpret most commands, including those used less frequently.
@@
During practice sessions, participants generated an average of 32 unique commands per task. The system successfully interpreted command variations, including different verbs (e.g., {grab, pick up, fetch, hold} the peach), object descriptions (e.g., peach, pink fruit), spatial references (e.g., first fruit, second watermelon from the left), and sentence structures (e.g., can you use the minimize button, peach inside the basket). Moreover, the system demonstrated the ability to interpret and decompose high-level commands and generalize beyond literal meanings. For example, it interpreted “increase the brightness” as (i) grab the brightness knob and (ii) twist right. In the example of “fold the window,” the system was able to connect fold to minimize, and press the minimize button.
@@
After practice sessions, as shown in Figure 7, participants continued to use diverse commands rather than relying solely on the most straightforward ones, and the system effectively handled this variability. Given that participants were free to use their own words without predefined commands or formats, these results highlight HandProxy’s capability to interpret flexible and varied input.
Figure 7. Command diversity in test sessions: Each task was repeated on three objects — one for practice and two for test sessions (top row: second object, bottom row: third object). We report the number of unique command phrasings (#), execution accuracy (Rate), normalized entropy ($H_{N}$), and a histogram showing the distribution of unique phrasings, sorted by frequency. We color-coded the successful attempts as blue and unsuccessful attempts as red to illustrate the distribution of correct/incorrect commands. Finally, we provide examples of successful and unsuccessful participant commands.
[…]
@@
To better understand HandProxy’s limitation, we analyzed the retried commands in the test sessions. In total, the system failed to execute or incorrectly executed 64 commands. Among them, 40 valid commands were categorized as invalid, and the system prompted users to try again and did not perform any action. Most were due to challenges for the LLM in identifying target objects (n=26), including synonyms (9), visual descriptions (6), object functionality (6), positional constraints (3), or other system errors (2). For example, participants described the power button as “the white and round button on the right side of the screen” (P1), the basket as “the brown object” (P16), or directly described the visual content within a widget, such as describing the window as “the dog photo” (P2). In other cases, participants referred to objects with synonyms or their expected functionality, such as describing the power button as “start on off button” (P10), or the heart button as “favorite button” (P16). While in many cases the system could infer beyond the object metadata, it struggled to account for the possible descriptions described above. This shows the need for a more comprehensive understanding of the objects and environment to support more diverse descriptive commands, which we will discuss in section 7.
@@
Apart from invalid commands with no action, the system incorrectly executed 24 valid commands. Among them, 3 were due to speech recognition errors, while 18 resulted from misinterpreted hand interactions. These include ambiguous commands (6), incorrect command decomposition (5), incorrect gesture parsing (4), or performed on the wrong target objects (3). Some commands had multiple possible interpretations, where context could have clarified the intended meaning. For instance, when P6 said, “turn the volume slider to maximum,” the participant expected the slider to move to the top. However, the LLM interpreted the word “turn” literally and the system did a “twist right” gesture instead. Errors also occurred when participants combined multiple commands into one sentence. For example, P20 said, “pinch the resize button and pull it to the right”, expecting a continuous pinch-and-move gesture. Instead, the LLM parsed “pull” as a separate command, initiated a “grab” gesture that canceled the pinch and caused an unexpected action. The system identified the wrong object from some other commands. For example, P5 said “pick up the first watermelon on the right” to refer to the rightmost watermelon, but the system interpreted it as the first watermelon overall (left to right).
@@
Nonetheless, participants were able to complete 100% of the tasks and recover from errors using alternative commands, with an average of 1.09 attempts (std 0.33) per command.
@@ 6.3. RQ3: What strategies, preferences, and methods do participants use for specifying intent?
Participants demonstrated diverse strategies and preferences while prompting the system. In practice sessions, the average command length was 5.25 words (std: 2.72, median: 5). In test sessions, the average was slightly lower at 4.73 words (std: 2.32, median: 4). For free exploration sessions – where users were not restricted to specific tasks – the average increased to 5.95 words (std: 3.32, median: 5). Participants who preferred more detailed commands found it easier to “specify what I want” (P2) or believed it would “help the system understand” their intent more accurately (P7). For example, to grab the peach, P15 used the command, “can you grab the peach and hold it in your hand?” In the interviews, 14 out of 20 participants favored shorter commands for their simplicity, clarity, and lower risk of errors. As P16 mentioned, “The longer it is, the more words I could say wrong and it could misinterpret.”
However, having shorter commands does not necessarily mean commands are always simple and low-level. In fact, participants used a mix of high-level and direct commands, such as “maximize the volume” (high-level) or “pinch the volume slider” followed by “pull up” (direct). For example, P9 mentioned that “I just wanted to tell the system what I want to do, and let the system figure out what gestures to do,” while others mentioned that using detailed control would be “more accurate” (P8), especially in precise tasks (P11) or those that need fine-grained controls (P6). Participants also reported the need to switch to a new mental model while using the system. This led some participants to use more descriptive ways of specifying their interaction intentions, such as “touch your index finger and thumb” to describe the pinch gesture (P13). As P17 said,
@@ 6.4. RQ4: How do participants experience and perceive HandProxy?
We evaluated system usability using 7-point Likert scale questions, covering system effectiveness, interaction variety, ease of use, responsiveness, learnability, consistency, and user confidence. The questions are based on the System Usability Scale (Brooke, 1996), with modifications to better align with the specific tasks in this user study. The Likert scale questions were administered three times at different stages to identify potential learning effects. However, no significant differences were found between them. Therefore, we report the results from the final administration that were conducted after all tasks were completed, as they provide the most comprehensive reflection of participants’ overall experience with the system.
Overall, participants found HandProxy effective for hand interactions (avg. 5.5, std. 0.82), can support various interaction gestures (average 5.75, std. 0.91), easy to use (average 5.7, std. 1.12), and participants were generally confident in using it (average 5.3, std. 1.08). Regarding system responsiveness, participants gave an average of 5.15 (std. 1.18). While participants appreciated the fast timing controls (e.g., stop, continue, undo) (e.g., P6), overall responsiveness across different types of commands could be improved for smoother interaction (P2, P8). HandProxy was also seen as intuitive and easy to learn (avg. 5.7, std. 1.41), especially given that the study was intentionally designed to have only a very brief on-boarding session. As P3 mentioned, “a few minutes of play and you should be ready.” However, P15 highlighted challenges for users unfamiliar with AR/VR, often relying on long, descriptive commands and struggling with terminology. Both P15 and P11 suggested that a tutorial with example commands, words, and objects would make the system more intuitive for beginners.
Additionally, participants valued HandProxy’s ability to handle high-level commands (P1, P2, P8, P11, P19); robustness to certain ambiguity in the command (P5, P10, P11); consistency on the same task (P7, P15); its memory capability (P19); and its ability to take continuous speech input and perform actions on the go (P7). For example, P1 mentioned that HandProxy is able to directly execute her high-level commands step by step, “even though I haven’t told the system how to perform the actual task.”
Figure 8. Final Likert scale responses of system usability.
@@
A figure of the final Likert scale responses across seven evaluation criteria related to the system’s usability and performance. The categories on the vertical axis include: Effectiveness of Initiating Hand Interactions, Variety of Interactions User Can Perform, Easiness of Use, Responsiveness of the System, Easiness to Learn, Consistency of the System Performance, Confidence of Using the System for Interaction Tasks. The majority of responses lean towards agreement and strong agreement, particularly in categories of ”Easiness of Use,” ”Easiness to learn,” ”System consistency,” and ”System variety”. Specifically, for Effectiveness of initiating hand interactions, majority agreed (55%), strongly agreeing (5%) or somewhat agreed (25%), with 15% neutral. For Variety of interactions user can perform, 60% agreed, 15% strongly agreed, 10% somewhat agreed, 15% neutral. For Ease of use, 50% agreed, 20% strongly agreed, 20% somewhat agreed, 10% somewhat disagreed. For responsiveness of the system, 40% somewhat agreed, 30% agreed, 10% strongly agreed, 15% somewhat disagreed, 5% neutral. For Ease to learn – 45% agreed, 30% strongly agreed, 10% neutral, and 5% each for somewhat agree, somewhat disagree, and disagree. For Consistency of the system performance, 30% agreed, 25% somewhat agreed, 20% strongly agreed, 15% neutral, 10% somewhat disagreed. For Confidence of using the system for interaction tasks, 50% agreed, 25% somewhat agreed, 5% strongly agreed, 10% each for neutral and somewhat disagreed.
@@ 6.5. RQ5: What other expectations do participants have for the system?
In the free exploration session, many participants tried commands requiring multimodal understanding of the environment, context, and how different objects are related to each other. For example, P18 wanted to say “like the dog” to press the heart button on top of the image viewer window. This expectation of environment understanding also extends to how the physics work in the virtual environment, such as “throw the apple to the dog” and “catch the apple right before it hits the table.” Commands based on visual descriptions were also used. As P19 mentioned,
@@
Participants also tried additional gestures that the system did not yet directly support, such as roll, throw, flip, wipe, and detailed finger controls. While there was no direct match to these gestures, HandProxy still tried to utilize existing gestures to reproduce the command to its best. For example, the system performed a grab and rotate to “flip a basket.” In addition, participants also suggested other hand controls that could be useful, including bimanual control (P1, P2, P13), supporting more precise, affordance-dependent object interaction (P2, e.g., grab the basket handle/side), and detailed individual finger controls (P5). Additionally, participants would like the system to understand object states and conditions in the environment. For example, P20 tried “if the photo has been resized, press the confirm button,” This also includes monitoring the state of the object while being manipulated by the virtual hand. For example, P13 mentioned that, “There was one time when it was trying to move the apple but didn’t realize it no longer had the apple in its hand,” if the system was aware of the hand state, it could automatically correct this error. These require understanding and monitoring the object state to relieve users from “monitoring conditions constantly.” (P19) Participants also emphasized the importance of additional feedback to help users identify and recover from errors. For example, P6 recommended underlining the parts in user’s command that are confusing or uninterpretable, enabling users to rephrase their input more effectively in subsequent attempts. P1 suggested having the system predict and suggest possible next moves after the command, such as showing all afforded hand interactions once the user grabbed an object. Additionally, participants noted that improving the speech recognition system and reducing overall latency could enhance the system’s usability. Current limitations and possible extentions on HandProxy are further discussed in section 7.